# Architectural Patterns for Designing Quantum Artificial Intelligence Systems

### v1

@@ Abstract
Utilising quantum computing technology to enhance artificial intelligence systems is expected to improve training and inference times, increase robustness against noise and adversarial attacks, and reduce the number of parameters without compromising accuracy. However, moving beyond proof-of-concept or simulations to develop practical applications of these systems while ensuring high software quality faces significant challenges due to the limitations of quantum hardware and the underdeveloped knowledge base in software engineering for such systems. In this work, we have conducted a systematic mapping study to identify the challenges and solutions associated with the software architecture of quantum-enhanced artificial intelligence systems. Our review uncovered several architectural patterns that describe how quantum components can be integrated into inference engines, as well as middleware patterns that facilitate communication between classical and quantum components. These insights have been compiled into a catalog of architectural patterns. Each pattern realises a trade-off between efficiency, scalability, trainability, simplicity, portability and deployability, and other software quality attributes.
@@
[inst3]organization=University of New South Wales,city=Sydney, postcode=2052, state=NSW, country=Australia
@@
[inst4]organization=Australian National University,city=Canberra, postcode=2601, state=ACT, country=Australia
@@
[inst5]organization=School of Physics, The University of Melbourne,city=Melbourne, postcode=3010, state=VIC, country=Australia
@@ 1 Introduction
Numerous studies have explored the physical, mathematical, and algorithmic principles underlying QML; however, only a few works have focused on application designs for existing quantum computers. Specifically, these applications are tailored to Noisy Intermediate-Scale Quantum (NISQ) devices, which are quantum processors with limited qubits and susceptibility to noise, that are available nowadays Preskill (2018). Despite these constraints, NISQ devices can offer a valuable platform for developing and testing quantum algorithms in AI applications Bharti et al. (2022); Coyle (2022), addressing various machine learning tasks such as classification and clustering. The inherent limitations of NISQ devices, coupled with non-functional requirements, such as performance, reliability, security, and usability, defined by software quality attributes Bi et al. (2021); ISO/IEC 25010:2023 , pose significant challenges in translating quantum algorithms into deployable and high-quality real-world applications.
@@ 2 Research Methodology
Figure 1 illustrates our evidence-based research methodology for identifying architectural patterns. The methodology is based on a systematic mapping study and is adhering to the guidelines outlined in Keele et al. (2007). The framework consists of two main steps: pattern identification and pattern classification. The pattern identification step encompasses several substeps, including paper search, paper collection, snowballing, quality assessment, and paper analysis. We specifically target high-quality academic papers related to architectural patterns in quantum computing systems from electronic databases such as the ACM Digital Library, [^1] https://dl.acm.org/ IEEE Xplore, [^2] https://ieeexplore.ieee.org/Xplore/home.jsp SpringerLink, [^3] https://link.springer.com/ Nature Portfolio, [^4] https://www.nature.com/nature-portfolio and Google Scholar. [^5] https://scholar.google.com/ The search string for all databases was composed of two parts, one related to quantum AI and the other to software engineering:
@@
3 Role of Architectural Patterns in Quantum AI Systems
Each pattern describes a high-level solution to a challenge encountered when designing, implementing, and/or managing AI systems Lu et al. (2023a). In this work, similar to Lu et al. (2024a), each pattern is formally represented using a template that includes the following components: name (with an optional graphical representation), summary, description of the associated problem, solution provided by the pattern, benefits, drawbacks, and known uses. In this work, we classify architectural patterns specific to quantum AI systems into two categories: quantum-classical split patterns and quantum middleware patterns.
Figure 2: Reference architecture for quantum-enhanced AI systems, featuring data pre-processing (Data PP), classical components (CC), and quantum components (QP), annotated with relevant architectural patterns.
@@ 3 Role of Architectural Patterns in Quantum AI Systems
System layer. This layer of the quantum AI system architecture is responsible for executing the trained models and making predictions or decisions based on input data. At this layer, architectural patterns are provided to integrate quantum and classical components, ensuring modularity, scalability, and efficient collaboration between various layers. The quantum components can be represented by quantum circuits, with design solutions illustrated in Figure 3. In this work, we have identified the following architectural patterns:
@@ 4.1 Quantum-classic split architectural patterns (SP)
All quantum-classical split patterns are defined within the context of hybrid classical-quantum AI systems. share the same context described here. In quantum computing, the hybrid quantum-classic algorithm represents a specific design solution, formalised as the design pattern known as the quantum-classic split Leymann (2019); Weigold et al. (2021b). Within quantum AI systems, real-world applications typically process classical data and produce classical outputs with few exceptions. One of such exceptions is simulating quantum objects, such as molecules and electronic systems Smith et al. (2019); Li et al. (2021a); Cheng et al. (2020). Additionally, some experimental setups operate exclusively on quantum data, performing classification of quantum states, which is crucial for validating quantum principles and exploring new algorithms Gao et al. (2022); Zoufal et al. (2019); Grant et al. (2018). For practical AI software applications, a hybrid quantum-classical software architecture is essential and, arguably, necessary. Our systematic mapping study confirms this, as all 113 selected papers on quantum AI systems discuss hybrid quantum-classical software architectures. In all cases, the extent of work delegated to the classical part versus the quantum part is a critical system-level architectural decision. Therefore, this work treats the quantum-classic split as a family of architectural patterns rather than a single design pattern. It is important to note that the quantum-classic split constitutes a unique family of architectural solutions exclusive to quantum software, with no equivalents in classical software engineering.
@@
We have identified seven architectural patterns related to the quantum-classical split, as compiled in Table 2. These patterns address the integration of quantum components within a classical inference engine in the field of quantum AI. The table provides the proposed names of these patterns, their summaries, and a comprehensive list of references to primary sources, capturing all known instances of their use. A comprehensive description of each identified quantum-classic split pattern is provided below, along with three representative use cases for each pattern.
@@
Problem: The problem to solve is enhancing expressibility of the quantum-enhanced model and increasing the number of trainable parameters within the constraints imposed by NISQ-era quantum computers, which limit the number of qubits and the depth of quantum circuits.
@@
    Jahin et al. (2023) [S52] and Schetakis et al. (2022) [S93] tested a hybrid architecture, which consists of an input layer, several classical neural network hidden layers, and one quantum layer. In these works, the authors applied this architecture to supply chain backorder prediction and abstract binary classification, respectively. To compute the gradients of quantum layers, the authors utilised PennyLane (see Bergholm et al. (2022)), which introduced the concept of differentiable quantum nodes that can be used in conjunction with classical layers.
@@
    Thakkar et al. (2024) used the intermediate quantum layer to improve financial forecasting.
@@
    Sebastianelli et al. (2022) and Sunkel et al Sünkel et al. (2023) applied this pattern in a hybrid architecture for multiclass classification of remote sensing imagery and classification of computed tomography scans of lungs to detect COVID-19, respectively.
@@
    Sakhnenko et al. (2022) used the intermediate quantum layer in the architecture of an autoencoder for anomaly detection in a dataset.
@@ 4.2 Quantum middleware architectural patterns (MP)
Service or micro-service wrapper (MP-1) In the context of quantum middleware, this pattern enables seamless communication and data exchange between classical and quantum components by wrapping quantum resources into classical services or small, loosely coupled, and independently deployable micro-services. [S20; S45; S53; S55; S80-81; S83; S92; S109]
@@
Quantum workflows orchestration (MP-3) In the context of hybrid quantum-classical workflows or quantum micro-service architecture, this pattern is designed to automatically deploy quantum components to the most suitable quantum provider, as well as provide automated scheduling of quantum programs to run continuously, or according to time-based or event-based schedules. [S2; S32; S47; S120-122]
@@
    Scalability and composability. This pattern enhances scalability and composability by leveraging specialised, unified, and standardised API functions, frequently based on OpenMPI frameworks [S20; S45; S83]. This approach ensures quantum services’ consistent and efficient scaling across various platforms.
@@
Summary: In the context of hybrid quantum-classical workflows Weder et al. (2020b, 2021a); Leymann and Barzen (2021) or quantum micro-service architecture O’Meara et al. (2023), this pattern is designed to automatically deploy quantum components to the most suitable quantum provider [S32, S82], as well as provide automated scheduling of quantum programs to run continuously, or according to time-based or event-based schedules [S109].
@@
Solution: To address these challenges, the software development and architectural designs of such systems necessitate tools that can structure software transactions into workflows. A workflow is defined as a technology that specifies the partial order of a collection of heterogeneous activities aimed at achieving a composite goal Leymann and Barzen (2021). The specific sequence of these activities is dictated by a workflow model. Typically represented as nodes in a directed graph, workflow models delineate the required tasks, their execution sequence, and the data flow between them. A workflow engine, or orchestrator, facilitates the automatic orchestration, deployment, and execution Weder et al. (2020b). The orchestrator is often implemented as middleware that analyses high-level programming language code, identifies quantum components, executes a classic-quantum split, and generates corresponding API functions. This middleware pattern, which can be implemented on either the server side or client side, generally includes components such as an orchestrator De Stefano et al. (2022); O’Meara et al. (2023) S109 S82, a language translator/generator De Stefano et al. (2022); Romero-Alvarez et al. (2024), and a scheduler De Stefano et al. (2022); O’Meara et al. (2023).
@@
    Simplicity. This pattern introduces a single point of failure, which means that if one part of the middleware fails, it could potentially disrupt the entire service, causing system-wide outages.
@@ 5 Rationale for using quantum computing in AI systems (RQ2)
Inference speed-up. Many papers report the potential for algorithmic inference speed-up in quantum AI through the use of quantum neural networks [S26; S77; S113; S125] and quantum kernels for support vector machines [S3; S98; S111]. This speed-up is typically associated with the reduced time complexity reported for quantum algorithms compared to their classical analogs [S26; S77; S98; S111; S113]. Several numerical experiments on classical simulations of quantum computers confirms faster execution times compared to classical algorithms for quantum neural networks [S36] and quantum kernels [S56]. In some cases, these advantages have been achieved even in simulations, laying the groundwork for quantum-inspired machine learning systems. Quantum-enhanced kernels for support vector machines have been also successfully run on an IBM quantum computer, although their execution times were not reported [S98; S111].
@@
Robustness against adversarial attacks. Robustness against well-engineered adversarial attacks is another potential advantage of quantum AI [S107]. With the development of quantum AI, researchers have raised concerns that inference engines could suffer from adversarial attacks similar to traditional neural networks [S25]. Indeed, it has been established in [S25] that quantum classifiers can be effectively deceived by minor adversarial perturbations. However, an adversarial training process, proposed by the authors, can significantly enhance their robustness against such attacks. The robustness of quantum neural networks against adversarial attacks has been confirmed in the context of fault classification for industrial control systems [S29] and image classification [S4, S8]. Additionally, it has been demonstrated that a quantum boost algorithm model trained with a quantum annealer is more robust to adversarial attacks compared to a state-of-the-art random forest model [S58]. General principles, perspectives, and challenges of quantum adversarial machine learning are discussed in [S107]. In this work, the authors have established that the adversarial vulnerability of quantum classifiers originates from their underlying Hilbert spaces. They have noted several defense strategies for quantum classifiers, including adversarial robustness through quantum noise, certifiable robustness, and adversarial training. Finally in [S8], West et al. (2023a) have benchmarked the robustness of several quantum variational classifiers and classical classifiers against adversarial attacks. The results show that quantum classifiers offer a notably enhanced robustness against classical adversarial attacks by learning features, which are not detected by the classical neural networks, indicating a possible quantum advantage.
Quantum sampling supremacy. Quantum mechanical systems can produce probability distributions that exhibit quantum correlations, which are challenging to capture using classical models [S22]. Quantum ML using generative models based on quantum neural networks facilitates efficient learning and loading of generic probability distributions with an exponential speed-up compared to classical algorithms [S22; S71; S100]. While this capability may not have immediate practical applications at the model level, it serves as a compelling demonstration of “quantum learning supremacy,” which holds significant interest for the advancement of quantum AI [S71; S100].

### v2

@@ Abstract
Utilising quantum computing technology to enhance artificial intelligence systems is expected to improve training and inference times, increase robustness against noise and adversarial attacks, and reduce the number of parameters without compromising accuracy. However, moving beyond proof-of-concept or simulations to develop practical applications of these systems while ensuring high software quality faces significant challenges due to the limitations of quantum hardware and the underdeveloped knowledge base in software engineering for such systems. In this work, we have conducted a systematic mapping study to identify the challenges and solutions associated with the software architecture of quantum-enhanced artificial intelligence systems. The results of the systematic mapping study reveal several architectural patterns that describe how quantum components can be integrated into inference engines, as well as middleware patterns that facilitate communication between classical and quantum components. Each pattern realises a trade-off between various software quality attributes, such as efficiency, scalability, trainability, simplicity, portability, and deployability. The outcomes of this work have been compiled into a catalogue of architectural patterns.
@@
[inst3]organization=University of New South Wales,addressline=, city=Sydney, postcode=2052, state=NSW, country=Australia
@@
[inst4]organization=Australian National University,addressline=, city=Canberra, postcode=2601, state=ACT, country=Australia
@@
[inst5]organization=School of Physics, The University of Melbourne,addressline=, city=Melbourne, postcode=3010, state=VIC, country=Australia
{highlights}
The unique software architectural solutions in quantum AI primarily focus on determining which tasks to delegate to a quantum computer (the quantum-classic split) and how quantum and classical software components communicate (interface and middleware).
This work identified existing architectural patterns that focus on integrating quantum components into inference engines and organising the interaction between classical components, quantum components, and hardware.
We anticipate that the pattern catalogue developed in this work will serve as guidance for software architects to effectively utilise quantum components in their projects.
Quantum AI systems offer several benefits, such as accelerating the inference and training stages, enhancing robustness against noise in data and adversarial attacks, and achieving better accuracy. However, most of the collected evidence confirming these advantages is obtained from simulations on classical computers.
@@ 1 Introduction
Numerous studies have explored the physical, mathematical, and algorithmic principles underlying QML; however, only a few works have focused on application designs for existing quantum computers. Specifically, these applications are tailored to Noisy Intermediate-Scale Quantum (NISQ) devices, which are quantum processors with limited qubits and susceptibility to noise, that are available nowadays Preskill (2018). Despite these constraints, NISQ devices can still offer a valuable platform for developing and testing quantum algorithms in AI applications Bharti et al. (2022); Coyle (2022), addressing various machine learning tasks such as classification and clustering. The inherent limitations of NISQ devices, coupled with non-functional requirements, such as performance, reliability, security, and usability, defined by software quality attributes Bi et al. (2021); ISO/IEC 25010:2023 , pose significant challenges in translating quantum algorithms into deployable and high-quality real-world applications.
@@ 2 Research Methodology
Figure 1 illustrates our evidence-based research methodology for identifying architectural patterns. The methodology is based on a systematic mapping study and is adhering to the guidelines outlined in Petersen et al. (2015). The framework consists of two main steps: pattern identification and pattern classification. The pattern identification step encompasses several substeps, including paper search, paper collection, snowballing, quality assessment, and paper analysis. We specifically target high-quality academic papers related to architectural patterns in quantum computing systems from electronic databases such as the ACM Digital Library, [^1] https://dl.acm.org/ IEEE Xplore, [^2] https://ieeexplore.ieee.org/Xplore/home.jsp SpringerLink, [^3] https://link.springer.com/ Nature Portfolio, [^4] https://www.nature.com/nature-portfolio and Google Scholar. [^5] https://scholar.google.com/ The search string for all databases was composed of two parts, one related to quantum AI and the other to software engineering:
@@
3 Reference architecture, role of architectural patterns in quantum AI and related works
Figure 2: Reference architecture for quantum-enhanced AI systems, featuring data pre-processing (Data PP), classical components (CC), and quantum components (QC), annotated with relevant architectural patterns.
@@ 3 Reference architecture, role of architectural patterns in quantum AI and related works
System layer. This layer of the quantum AI system architecture is responsible for executing the trained models and making predictions or decisions based on input data. The architectural patterns identified in this work belong to this layer and offer solutions to the challenges associated with integrating quantum and classical components, ensuring modularity, scalability, and efficient collaboration between various layers. The quantum components can be represented by quantum circuits, with design solutions illustrated in Figure 3. In this work, we have identified the following architectural patterns, described in detail in Section 4:
@@
Note that several publications have already addressed the identification of architectural and design patterns in quantum software Leymann (2019); Weigold et al. (2021b, a); Georg et al. (2023); Khan et al. (2022), including the proposal of a specialised pattern language Leymann (2019). While these works focus on quantum software in general, this paper specifically concentrates on architectural patterns relevant to quantum AI and quantum machine learning.
@@ 4 Pattern catalog for quantum AI systems (RQ1)
This section provides a detailed description of the architectural patterns identified during the systematic mapping study. Each pattern describes a high-level solution to a challenge encountered when designing, implementing, and/or managing AI systems Lu et al. (2023a). In this work, similar to Lu et al. (2024a), each pattern is formally represented using a template that includes the following components: name (with an optional graphical representation), summary, description of the associated problem, solution provided by the pattern, benefits, drawbacks, and known uses. In this work, we classify architectural patterns specific to quantum AI systems into two categories: quantum-classical split patterns (subsection 4.1) and quantum middleware patterns (subsection 4.2) .
@@ 4.1 Quantum-classic split architectural patterns (SP)
All quantum-classical split patterns are defined within the context of hybrid classical-quantum AI systems. In quantum computing, the hybrid quantum-classic algorithm represents a specific design solution, formalised as the design pattern known as the quantum-classic split Leymann (2019); Weigold et al. (2021b). Within quantum AI systems, real-world applications typically process classical data and produce classical outputs with few exceptions. One of such exceptions is simulating quantum objects, such as molecules and electronic systems Smith et al. (2019); Li et al. (2021a); Cheng et al. (2020). Additionally, some experimental setups operate exclusively on quantum data, performing classification of quantum states, which is crucial for validating quantum principles and exploring new algorithms Gao et al. (2022); Zoufal et al. (2019); Grant et al. (2018). For practical AI software applications, a hybrid quantum-classical software architecture is essential and, arguably, necessary. Our systematic mapping study confirms this, as all 113 selected papers on quantum AI systems discuss hybrid quantum-classical software architectures. In all cases, the extent of work delegated to the classical part versus the quantum part is a critical system-level architectural decision. Therefore, this work treats the quantum-classic split as a family of architectural patterns rather than a single design pattern. It is important to note that the quantum-classic split constitutes a unique family of architectural solutions exclusive to quantum software, with no equivalents in classical software engineering.
@@
We have identified seven architectural patterns related to the quantum-classical split, as compiled in Table 2. These patterns address the integration of quantum components within a classical inference engine in the field of quantum AI. The table provides the proposed names of these patterns, their summaries, and a complete list of primary sources, capturing all known instances of their use. A comprehensive description of each identified quantum-classic split pattern is provided below, along with three representative use cases for each pattern.
@@
Problem: The problem to solve is enhancing expressiveness of the quantum-enhanced model and increasing the number of trainable parameters within the constraints imposed by NISQ-era quantum computers, which limit the number of qubits and the depth of quantum circuits.
@@
    Jahin et al. (2023) [S52] and Schetakis et al. (2022) [S93] tested a hybrid architecture, which consists of an input layer, several classical neural network hidden layers, and one quantum layer. In these works, the authors applied this architecture to supply chain backorder prediction and abstract binary classification, respectively. To compute the gradients of quantum layers, the authors in Bergholm et al. (2022) utilised PennyLane, which introduced the concept of differentiable quantum nodes that can be used in conjunction with classical layers.
@@
    Thakkar et al. (2024) [S39] used the intermediate quantum layer to improve financial forecasting.
@@
    Sebastianelli et al. (2022) [S49] and Sünkel et al. (2023) [S33] applied this pattern in a hybrid architecture for multiclass classification of remote sensing imagery and classification of computed tomography scans of lungs to detect COVID-19, respectively.
@@
    Sakhnenko et al. (2022) [S30] used the intermediate quantum layer in the architecture of an autoencoder for anomaly detection in a dataset.
@@ 4.2 Quantum middleware architectural patterns (MP)
Service or micro-service wrapper (MP-1) In the context of quantum middleware, this pattern enables seamless communication and data exchange between classical and quantum components by wrapping quantum resources into classical services or small, loosely coupled, and independently deployable micro-services. [S20; S44-45; S53; S55; S80-81; S83; S92; S96; S109]
@@
Quantum workflows orchestration (MP-3) In the context of hybrid quantum-classical workflows or quantum micro-service architecture, this pattern is designed to automatically deploy quantum components to the most suitable quantum provider, as well as provide automated scheduling of quantum programs to run continuously, or according to time-based or event-based schedules. [S2; S32; S42; S47; S82]
@@
Note that the quantum software full-stack often follows a layered architecture, as presented in [S44]. When applied within the context of a service-oriented architecture, this layered structure gives rise to a service layer pattern. In quantum systems, the service layer functions as a wrapper for higher-level abstractions rather than organising business logic, making it functionally equivalent to the service wrapper pattern.
@@
    Scalability and composability. This pattern enhances scalability and composability by leveraging specialised, unified, and standardised API functions [S20; S45; S83].
@@
Summary: In the context of hybrid quantum-classical workflows Weder et al. (2020, 2021a); Leymann and Barzen (2021) or quantum micro-service architecture O’Meara et al. (2023), this pattern is designed to automatically deploy quantum components to the most suitable quantum provider [S32, S82], as well as provide automated scheduling of quantum programs to run continuously, or according to time-based or event-based schedules [S109].
@@
Solution: To address these challenges, the software development and architectural designs of such systems necessitate tools that can structure software transactions into workflows. A workflow is defined as a technology that specifies the partial order of a collection of heterogeneous activities aimed at achieving a composite goal Leymann and Barzen (2021). The specific sequence of these activities is dictated by a workflow model. Typically represented as nodes in a directed graph, workflow models delineate the required tasks, their execution sequence, and the data flow between them. A workflow engine, or orchestrator, facilitates the automatic orchestration, deployment, and execution Weder et al. (2020). The orchestrator is often implemented as middleware that analyses high-level programming language code, identifies quantum components, executes a classic-quantum split, and generates corresponding API functions. This middleware pattern, which can be implemented on either the server side or client side, generally includes components such as an orchestrator De Stefano et al. (2022); O’Meara et al. (2023) [S109, S82], a language translator/generator De Stefano et al. (2022); Romero-Alvarez et al. (2024), and a scheduler De Stefano et al. (2022); O’Meara et al. (2023).
@@
    Reliability. This pattern introduces a single point of failure, which means that if one part of the middleware fails, it could potentially disrupt the entire service, causing system-wide outages.
@@ 5 Rationale for using quantum computing in AI systems (RQ2)
Inference speed-up. Many papers report the potential for algorithmic inference speed-up in quantum AI through the use of quantum neural networks [S26; S77; S113; S125] and quantum kernels for support vector machines [S3; S98; S111]. This speed-up is typically associated with the reduced time complexity reported for quantum algorithms compared to their classical analogs [S26; S77; S98; S111; S113]. Note that, given the slow operational speed of existing quantum gates, transferring the speedup associated with reduced time complexity in algorithms to actual speedup on quantum hardware, relative to classical systems, may require problem sizes that are unrealistically large. Several numerical experiments on classical simulations of quantum computers confirms faster execution times compared to classical algorithms for quantum neural networks [S36] and quantum kernels [S56]. The fact that speed-up can be achieved even in simulations lays the foundation for quantum-inspired machine learning systems. Quantum-enhanced kernels for support vector machines have been also successfully run on an IBM quantum computer, although their execution times were not reported [S98; S111].
@@
Robustness against adversarial attacks. Robustness against well-engineered adversarial attacks is another potential advantage of quantum AI [S107]. With the development of quantum AI, researchers have raised concerns that inference engines could suffer from adversarial attacks similar to traditional neural networks [S25]. Indeed, it has been established in [S25] that quantum classifiers can be effectively deceived by minor adversarial perturbations. However, an adversarial training process, proposed by the authors, can significantly enhance their robustness against such attacks. The robustness of quantum neural networks against adversarial attacks has been confirmed in the context of fault classification for industrial control systems [S29] and image classification [S4, S8]. Additionally, it has been demonstrated that a quantum boost algorithm model trained with a quantum annealer is more robust to adversarial attacks compared to a state-of-the-art random forest model [S58]. General principles, perspectives, and challenges of quantum adversarial machine learning have been discussed in [S107]. In this work, the authors have established that the adversarial vulnerability of quantum classifiers originates from their underlying Hilbert spaces. They have noted several defence strategies for quantum classifiers, including adversarial robustness through quantum noise, certifiable robustness, and adversarial training. Finally in [S8], West et al. (2023a) have benchmarked the robustness of several quantum variational classifiers and classical classifiers against adversarial attacks. The results show that quantum classifiers offer a notably enhanced robustness against classical adversarial attacks by learning features, which are not detected by the classical neural networks, indicating a possible quantum advantage.
Quantum sampling supremacy. Quantum mechanical systems can produce probability distributions that exhibit quantum correlations, which are challenging to capture using classical models [S22]. Quantum generative models based on quantum neural networks facilitates efficient learning and loading of generic probability distributions with an exponential speed-up compared to classical algorithms [S22; S71; S100]. While this capability may not have immediate practical applications at the model level, it serves as a compelling demonstration of “quantum learning supremacy,” which holds significant interest for the advancement of quantum AI [S71; S100].